# Ultrafast optical switching to a heterochiral charge-density wave state

### v1

@@ Acknowledgments
Author contributions
The project was conceived by C.R.; W.C.H. conducted experiments and data analysis with contributions from Y.S.L., F.K., and J.G.H; the 1T-TaS2 samples were provided by K.R.; S.M., S.-S.H., and H.-T.J. carried out the DFT simulations with contributions from W.C.H.; G.v.W. performed the dynamical LEED simulations; results of DFT and dynamical LEED simulations were analyzed by W.C.H., S.-S.H., H.-T.J., G.v.W., and S.M.; the manuscript was written by W.C.H. and C.R. with contributions from G.v.W, J.G.H., and F.K.. All authors discussed the results and commented on the manuscript.
I Methods
1. LEED measurements In the experiments, large flakes of 1T-TaS2 samples (diameter 2-3 mm, thickness 150 $\upmu$m) are cleaved under high vacuum and transferred to an ultrahigh vacuum chamber with a base pressure of $2\times 10^{\textrm{-}10}$ mbar. The samples are prepared in the C-phase at a temperature of 30 K using a liquid helium flow cryostat. LEED images in a backscattering geometry near normal incidence (cf. Fig. 1b) are recorded using a narrow electron beam (150 $\upmu$m diameter) with high momentum resolution from a millimeter-sized photoelectron gun (sample-to-gun distance 10 mm, see Ref. Vogelgesang2018 for details on the electron source). The low-temperature C-phase of 1T-TaS2 exhibits a ($\sqrt{13}\times\sqrt{13}$)R13.9 CDW superstructure. As our study is concerned with the chirality of the superstructure Zong2018 ; Ishioka2010 ; Song2022 ; Liu2023 , in each measurement, we begin with a large single-crystalline area of a single CDW chirality. The diffraction pattern of the initial state (cf. Fig. 1d, top-left panel) exhibits a hexagonal arrangement with bright main lattice peaks surrounded by weaker CDW superstructure peaks, which arise from the CDW-coupled periodic lattice distortion Overhauser1971 ; Eichberger2010 .
2. Electron-laser spatial overlap We use single or double laser pulses to excite the sample under an angle of incidence of about 45 degrees. The laser spot size on the sample is 300 $\upmu$m (FWHM), ensuring uniform illumination of the area probed by the electron beam. The spatial overlap between the electron and laser beams is found using the technique of ultrafast low-energy electron diffraction (ULEED) Vogelgesang2018 . Specifically, at low-fluence excitation, we optimize the electron-laser overlap by maximizing the laser-induced Debye-Waller effect, i.e., the transient suppression of diffraction spot intensities.
3. Generation of single- and double-pulse The optical quench of 1T-TaS2 samples is conducted using either single or double laser pulses. Single pulses are isolated with a Pockels cell inside a laser amplifier (center wavelength 1030 nm). Femtosecond excitation pulses of about 200 fs duration (center wavelength 800 nm, FWHM spectral width 7 nm) are produced by an optical parametric amplifier. Double pulses with a pulse separation $|\tau|\leq 2\,\textrm{ps}$ are generated by passing a single pulse through a Michelson interferometer. A half-inch polarizing beamsplitter cube is used to split the pulse, and a quarter-wave plate is used in each arm of the interferometer to optimize the total laser power output. The two output pulses are orthogonally polarized with a measured interferometric visibility of 3 %.
4. Reversal from the heterochiral to the monochiral state The duration of the annealing laser pulse train is controlled by a mechanical shutter that produces a 1 s top-hat function with rising and falling edges spanning about 10 ms. Given sufficient fluence, the annealing laser pulse train can completely reverse the heterochiral state to a monochiral state (cf. Fig. 2c). This procedure is used for initializing each measurement. Incomplete reversal was observed if a sharper drop of the pulse train envelope was used.
5. Density-functional theory simulations Density functional theory (DFT) simulations are performed for the CDW moiré superstructure using projector augmented wave (PAW) Blochl1994 potentials as implemented in the Vienna Ab-initio Simulation Package (VASP) Kresse1993 ; Kresse1996 . A 350 eV kinetic energy cutoff in the plane-wave expansion is used. The PAW pseudopotentials correspond to the valence-electron configuration 5d36s2 for Ta and 3s23p4 for S. The exchange-correlation is treated within the generalized gradient approximation, as parameterized by Perdew, Burke, and Ernzerhof Perdew1996 . The Coulomb correlations within the 5d shells of Ta are described using the spherically averaged DFT+$U$ method Dudarev1998 , where the Hamiltonian only depends on the difference between on-site interaction $U$ and Hund’s exchange $J$, i.e., $U_{\textrm{eff}}=U-J$. We adopt the value of $U_{\textrm{eff}}=2.0$ eV for the Ta 5d orbitals Darancet2014 . Spin-orbit coupling is ignored in all simulations Ritschel2018 . For structural relaxation, the lattice constants $a=43.68\,\text{\AA}$ (13 times of the $1\times 1$ primitive cell lattice constant $3.36\,\text{\AA}$) and $c=11.8\,\text{\AA}$ (twice the Ta-Ta interlayer distance $5.9\,\text{\AA}$) are fixed vonWitte2019 ; Spijkerman1997 ; Givens1977 , while the positions of 1014 atoms are relaxed by a $\Gamma$-centered $1\times 1\times 2$ k-point mesh until the maximum force on each atom becomes less than 10 meV/Å. A $\Gamma$-centered $3\times 3\times 2$ k-point mesh is used for calculating the self-consistent ground state charge and electronic band structures. The total energies are converged to 10-${}^{\textrm{-}4}$ eV. In a separate calculation, The 1T-TaS2 C-phase is simulated by a 156-atom AL-stacking unit cell using a $\Gamma$-centered $6\times 6\times 6$ k-point mesh Ritschel2018 ; Lee2019 .

### v2

@@ Acknowledgments
Supplementary Information
This supplement provides detailed information about the experimental observation and theoretical modeling of the heterochiral charge-density wave state (H-state). These include spot profile analysis, dynamical LEED simulations, as well as electronic structure, atomic lattice, and charge texture analysis based on density-functional theory (DFT) simulations. Additional data sets of diffraction intensity as a function of diffraction spot position and the probing electron energy (LEED I-V curves) are also provided.
.2 I. Spot profile analysis
Figure S1 presents the lineouts of the main lattice, $\alpha$, and $\beta$ peaks (cf. main text Fig. 1e, right column). The lineout of the main lattice peak is along the horizontal direction, while the lineouts of the superstructure peaks are along the main-$\alpha$ or main-$\beta$ directions. In Fig. S1a, we find negligible change in the linewidths of the main lattice and $\alpha$ peaks after the optical quench. In Fig. S1b, lineouts of the main lattice, $\alpha$, and $\beta$ peaks in the H-state are compared, and their respective FWHMs are measured $\Delta k_{\text{main}}=0.28\,\textrm{nm}^{\textrm{-}1}$, $\Delta k_{\alpha}=0.40\,\textrm{nm}^{\textrm{-}1}$, and $\Delta k_{\beta}=0.38\,\textrm{nm}^{\textrm{-}1}$. The similar linewidths and lineouts of $\alpha$ and $\beta$ peaks evidence that the initial $\alpha$ chirality and the emergent $\beta$ chirality are both long-range order in the H-state. To estimate the lower bounds of $\alpha$ and $\beta$ domain sizes in the H-state, we approximate the diffraction lineouts in Fig. S1b as Gaussians and calculate the corresponding correlation lengths using the deconvolution formula $\xi_{\alpha,\beta}=2\pi/\sqrt{\Delta k_{\alpha,\beta}^{2}-\Delta k_{\text{main}}^{2}}$, where we assume that the instrument response function follows the lineout of the main lattice peak. The retrieved minimum correlation lengths are $\xi_{\alpha}=22$ nm and $\xi_{\beta}=24$ nm for $\alpha$ and $\beta$ domains in the H-state.
Figure S1: Lineouts of the main lattice, $\alpha$, and $\beta$ peaks. a, Lineouts of the main lattice and $\alpha$ peaks, compared between the C-phase and the H-state. As the sample switches from the C-phase to the H-state after the optical quench, the diffraction intensities of the main lattice and $\alpha$ peak are slightly reduced, but there is negligible change in the linewidth. b, Lineout comparison in the H-state. The lineout of the emergent $\beta$ peak is similar to that of the $\alpha$ peak. Lineout angles for the $\alpha$ and $\beta$ peaks are along the radial directions, i.e., the main-$\alpha$ and main-$\beta$ directions (cf. main text Fig. 1e, left column). The lineout angle for the main lattice peak is along the horizontal direction.
.3 II. Experimental LEED I-V curves and dynamical LEED simulations
In this section, we investigate how the H-state signal varies with the diffraction spot position and the probing electron energy (LEED I-V curves). Dynamical LEED simulations are performed for comparison.
The H-state signal is quantified by the normalized intensity change, $\Delta I_{\text{H-C}}/I_{\text{C}}\equiv(I_{\text{H}}-I_{\text{C}})/I_{\text{C}}$, where $I_{\text{H,C}}$ is the diffraction spot intensity in the H-state or the C-phase, normalized to the total count of the respective LEED image. In Fig. S2, the spot intensities of all $\alpha$ peaks decrease after the sample switches from the C-phase to the H-state. As a function of $\alpha$ spot position, $\Delta I_{\text{H-C}}/I_{\text{C}}$ fluctuates around -6.6 % with a standard deviation of 3.4 %. As a function of $\beta$ spot position, $\Delta I_{\text{H-C}}/I_{\text{C}}$ fluctuates around 75 % with a standard deviation of 36 % (Fig. S3). In both cases, the relative fluctuation is about 50 %. We note that $I_{\text{C}}$ of the $\beta$ peak takes the local background value, as $\beta$ peaks are absent in the C-phase. In Fig. S4, $I_{\text{H,C}}$ and $\Delta I_{\text{H-C}}/I_{\text{C}}$ show only a mild modulation as a function of the probing electron energy.
We investigate with dynamical LEED simulations how multiple scattering effects respond to a drastic structural change in the subsurface layer and compare the simulation results with experimental observations (Figs. S2S3, and S4). The simulation is based on our previous surface crystallographic analysis of the 1T-TaS2 C-phase [1], using the TensErLEED package [2]. We construct a two-layer C-phase structure and a hypothetical two-layer M-phase structure. The C-phase has the charge-density wave (CDW) induced periodic lattice distortion (PLD) on both surface and subsurface layers, while the M-phase has the surface layer with PLD and the subsurface layer without. By changing from the C-phase structure (surface: CDW, subsurface: CDW) to the M-phase structure (surface: CDW, subsurface: metallic), we simulate the effect of switching from the 1T-TaS2 C-phase (surface: $\alpha$ chirality, subsurface: $\alpha$ chirality) to the H-state (surface: $\alpha$ chirality, subsurface: $\beta$ chirality), assuming the subsurface $\beta$-layer texture in Fig. 3b of the main text. The simulation shows that a drastic structural change in the subsurface layer can result in large intensity fluctuations across all existing diffraction spots. Specifically, the spot intensity change $\Delta I_{\text{M-C}}/I_{\text{C}}$ takes both positive and negative values (Fig. S5d). This result is qualitatively different from our experimental observation in Fig. S4c, where all spot intensities of the $\alpha$ peak decrease after switching from the C-phase to the H-state.
Figure S2: Intensity of the $\alpha$ peak as a function of diffraction spot position. a, Experimental LEED image of the H-state (cf. main text Fig. 1d). Selected $\alpha$ peaks are labeled by spot indices a1 to a18. b, Normalized $\alpha$-peak intensity as a function of diffraction spot position, compared between the C-phase and the H-state. All spot intensities decrease after the sample switches to the H-state. c, Normalized intensity change of the $\alpha$ peak as a function of diffraction spot position. The average intensity change is -6.6 % (red dashed line).
Figure S3: Intensity of the $\beta$ peak as a function of diffraction spot position. a, Experimental LEED image of the H-state (cf. main text Fig. 1d). Selected $\beta$ peaks are labeled by spot indices b1 to b18. b, Normalized $\beta$-peak intensity as a function of diffraction spot position, compared between the C-phase and the H-state. All spot intensities increase after the sample switches to the H-state. c, Normalized intensity change of the $\beta$ peak as a function of diffraction spot position. The average intensity change (comparing to the background) is 75 % (red dashed line).
Figure S4: Experimental LEED I-V curves compared between the C-phase and the H-state. a, Experimental LEED image of the H-state. The H-state is generated with a single-pulse quench at a fluence of 1.6 mJ cm-2. The probing electron energy is 100 eV. Selected $\alpha$ and $\beta$ peaks are labeled by spot indices a1 to a14 and b1 to b14, respectively. b, Experimental LEED I-V curves for $\alpha$ and $\beta$ peaks in the C-phase and the H-state. c, Normalized intensity change of the $\alpha$ peak as a function of the probing electron energy and diffraction spot position. d, Normalized intensity change of the $\beta$ peak as a function of the probing electron energy and diffraction spot position.
Figure S5: Summary of the dynamical LEED simulations for two structures that differ in the subsurface layer. a, Simulated LEED pattern of the C-phase structure. The LEED pattern is a graphical representation of the result of the dynamical LEED simulation, assuming 2D Gaussian spot profiles with the peak height determined by the simulation. Normal incidence is assumed for the probing electrons in the simulation. The effective three-fold symmetry divides the main lattice and CDW peaks into two groups (yellow and green triangles). Distinct first-order CDW peaks are labeled by spot indices a1 to a14. b, Normalized intensity as a function of diffraction spot position, compared between the C- and M-phases. Due to multiple scattering effects, about half of the spots (a1, a2, a5, a6, a7, a9, a13, a14) have stronger intensity in the C-phase, while the other half (a3, a4, a8, a10, a11, a12) have stronger intensity in the M-phase. c, Simulated LEED I-V curves of the subsurface layer contribution. The normalized intensity change, defined as $\Delta I_{\text{M-C}}/I_{\text{C}}\equiv(I_{\text{M}}-I_{\text{C}})/I_{\text{C}}$, quantifies the contribution of the subsurface layer in the total CDW diffraction intensity. The diffraction intensity in the M- and C-phases, $I_{\text{M,C}}$, is averaged over all 14 CDW spots (a1 to a14) at each probing electron energy. The intensity fluctuation as a function of the probing electron energy is a result of the multiple scattering effect. At 100 eV, the subsurface layer contribution is 3.7 %. d, Simulated LEED I-V curves of the subsurface layer contribution at individual diffraction spots. Here, $I_{\text{M,C}}$ is the intensity of individual diffraction spots. As a function of the probing electron energy, $\Delta I_{\text{M-C}}/I_{\text{C}}$ fluctuates between positive and negative values for all spots. This result is qualitatively different from our experimental observation in Fig. S4c, where the intensity change takes only negative values.
.4 III. Mixed stacking order and the double-ring charge texture of the CDW moiré superstructure
The CDW moiré superstructure has the $\alpha$ chirality in one layer and the $\beta$ chirality in the other (Fig. S6a). The $27.8^{\circ}$ rotational mismatch between CDWs in the $\alpha$ and $\beta$ layers gives rise to a commensurate moiré supercell of 1014 atoms (black dashed hexagon in Fig. S6b). In each layer, there are 13 Ta hexagrams, forming a super-hexagram in a fractal-type arrangement (Fig. S6a). The angle between $\alpha$ and $\beta$ super-hexagrams gives rise to a mixed stacking order, realizing all 13 possible stacking types (Fig. S6c) within the CDW moiré unit cell. Specifically, the A-stacking locates at the center of the super-hexagram, the six I-, G-, L-, F-, H-, C-stackings coincide with the inner ring of the super-hexagram, and the six K-, B-, E-, D-, M-, J-stackings reside at the outer ring (cf. Fig. S6b).
In our DFT simulations, the atomic positions are relaxed, while the lattice constants of the CDW moiré unit cell, $a=43.68\,\text{\AA }$ (13 times of the $1\times 1$ primitive cell lattice constant $3.36\,\text{\AA }$) and $c=11.8\,\text{\AA }$ (twice the Ta-Ta interlayer distance $5.90\,\text{\AA }$), are fixed. Prior to structural relaxation, the 13 Ta hexagrams in each of the $\alpha$ and $\beta$ layer of the CDW moiré unit cell are equivalent and the electron density should distribute evenly among them. However, in the presence of an electronic reconstruction, which occurs as a result of structural transformation (cf. Fig. S7), the electron density redistributes into the inner and outer rings of the super-hexagrams, forming a double-ring charge texture.
To understand how the double-ring electron density distribution correlates with the atomic structure of the CDW moiré unit cell, we perform self-consistent energy calculations for the five nonequivalent stackings (A, B, K, H, C) of the standard 1T-TaS2 structure. The five stackings are representative of all 13 stacking types due to the three-fold symmetry of 1T-TaS2. The calculated electrostatic (Hartree) energy characterizes the Coulomb repulsion associated with each stacking type and may help to understand the stacking-dependent electron distributions in the CDW moiré superstructure.
[…]
Figure S6: Mixed stacking order of the CDW moiré superstructure. a, Schematics of Ta super-hexagrams in the $\alpha$ and $\beta$ layers. b, DFT simulated atomic structure of the light-induced CDW moiré superstructure in 1T-TaS2. Blue and red dots represent Ta atoms in the $\alpha$ and $\beta$ layers, respectively. Yellow dots represent S atoms. The $\alpha$ layer (blue) is on the top and the $\beta$ layer (red) is at the bottom. Here the moiré unit cell (black dashed hexagon) is obtained through the construction of a Wigner-Seitz cell with the lattice constants $a=43.68\,\text{\AA }$ and $c=11.8\,\text{\AA }$. For comparison, Ta hexagrams in the $\alpha$ and $\beta$ super-hexagram unit cells (cf. a) are highlighted by blue and red thick lines. The angle between the $\alpha$ and $\beta$ super-hexagrams give rise to a mixed stacking order. Namely, all 13 possible stacking types (denoted by capital letters A to M) are realized in the moiré unit cell. The A-stacking at the center has the highest Hartree energy. The I-, G-, L-, F-, H-, C-stackings at the inner ring (orange) of the super-hexagrams have the lowest Hartree energy. The K-, B-, E-, D-, M-, J-stackings at the outer ring (purple) of the super-hexagrams have the intermediate Hartree energy. c, Illustration of the 13 stacking types according to Ref. [3]. The capital letters mark the center of the Ta hexagram on the top of each stacking pair (i.e., Ta hexagrams in the $\alpha$ layer). Darker filling color corresponds to higher electrostatic (Hartree) energy.
Stacking type Hartree energy, E (eV per hexagram) Location in the super-hexagrams
A 0 center
B, D, J -8.027 outer ring
K, E, M -8.155 outer ring
H, I, L -15.317 inner ring
C, G, F -15.422 inner ring
Table S1: Electrostatic (Hartree) energy of standard 1T-TaS2 lattice structures. Equivalent stacking types are grouped in the 1st column. The Hartree energy of the A-stacking is 1055.407 eV per hexagram. Energies of other stackings are referenced to the A-stacking. The DFT simulations are performed over a $\Gamma$-centered $6\times 6\times 4$ k-point mesh. In the simulation, the CDW unit cell lattice constants ($a=12.1147\,\text{\AA }$, $c=11.8\,\text{\AA }$) are fixed, while the atomic positions are relaxed.
.5 IV. Emergent kagome structural order in the CDW moiré superstructure
The DFT simulation for the CDW moiré superstructure starts with a two-layer supercell, each layer mimicking the surface-layer CDW superstructure in the 1T-TaS2 C-phase [1]. The CDW superstructure has two chiral orientations and a rotation angle of $13.9^{\circ}$ with respect to the main lattice (cf. Fig. S6a). We denote $13.9^{\circ}$ (clockwise) as the $\alpha$ chirality and -$13.9^{\circ}$ (counter-clockwise) as the $\beta$ chirality. The unit cell of the CDW moiré superstructure is built with the $\alpha$ chirality in the top layer (layer 1) and the $\beta$ chirality in the bottom layer (layer 2).
In the DFT simulation, the lattice constants of the moiré unit cell ($a=43.68\,\text{\AA }$, $c=11.8\,\text{\AA }$) are fixed, while the atomic positions are relaxed. The relaxed atomic structure experiences a structural transformation from a $(\sqrt{13}\times\sqrt{13}$)R13.9 triangular superlattice to a $13\times 13$ kagome superlattice (Fig. S7b). The kagome superlattice features corner-sharing triangles with three nonequivalent sites (Fig. S7c). Comparison between the relaxed atomic structure and the C-phase structure (AL-stacking) gives insight to the emergence of the kagome structural order. Namely, displacements of the Ta hexagram centers (Fig. S8) break the translational symmetry of Ta hexagram in both $\alpha$ and $\beta$ layers, transforming the Ta superstructure in each layer from $(\sqrt{13}\times\sqrt{13}$)R13.9 to $13\times 13$.
Figure S7: Emergent kagome structural order in the CDW moiré superstructure. a, The $(\sqrt{13}\times\sqrt{13}$)R13.9 triangular lattice of the C-phase Ta superstructures. Only the Ta hexagram centers are shown (blue dots). Bonds of $12.1147\,\text{\AA }$ are plotted to illustrate the triangular lattice. b, The $13\times 13$ kagome lattice of the CDW moiré superstructure. The structural deformation caused by the coupling between $\alpha$ and $\beta$ layers clusters the 13 Ta hexagram centers into a super-hexagram. The super-hexagram has a three-fold symmetry, and bonds between neighboring super-hexagrams ($1182\,\text{\AA }-12.12$$1242\,\text{\AA }$) are longer than those within the super-hexagram ($12.0999\,\text{\AA }$-$12.1171\,\text{\AA }$). c, Closeup of the kagome superlattice with all atoms displayed. The kagome superlattice features corner-sharing triangles with three nonequivalent sites A, B, and C. Atoms near these three sites have orientations differing by $120^{\circ}$.
Figure S8: Comparison between Ta positions in the CDW moiré superstructure and the C-phase. a, Ta displacement from the C-phase structure in the $\alpha$ layer. (Top) The Ta displacements (depicted by arrows) can be grouped by 28 concentric rings according to the distance between the displaced Ta atom and the center of the $\alpha$ super-hexagram. As there are 169 Ta atoms within the super-hexagram unit cell, each ring is associated with 6 Ta atoms. The Ta atom at the center of the super-hexagram experiences no displacement. Ring #9 and #22 coincide with the inner and outer rings of the super-hexagram, where the Ta hexagram centers reside. (Bottom) Displacement amplitudes associated with each ring are shown by black dots, with the average value shown by red circles. The overall average of Ta displacement amplitude is around 0.04 Å  (red dashed line), which is about 20 % of the CDW distortion in the C-phase. b, Ta displacement from the C-phase structure in the $\beta$ layer. Displacements of Ta atoms in the $\alpha$ and $\beta$ layers have similar amplitudes but opposite sense of rotation with respect to the super-hexagram centers.
.6 V. Electronic band structures of the CDW moiré superstructure
Here we present the electronic band structures obtained from DFT simulations, using the Vienna Ab initio Simulation Package [4, 5, 6] with an on-site $U_{\textrm{eff}}$ of 2.0 eV for Ta 5d orbitals [7] in the generalized gradient approximation of the Perdew Burke Ernzerhof (PBE) form [8]. After relaxing the atomic positions (cf. Fig. S7), we perform self-consistent electronic structure calculations with a $3\times 3\times 2$ $\Gamma$-centered Monkhorst-Pack k-point mesh. The cutoff energy for the plane-wave basis is 400 eV. The band structure shown in Fig. S9a is calculated along paths connecting the 6 high-symmetry k-points in the 1st Brillouin zone of the CDW moiré unit cell. In stark contrast to the insulating band structure of the C-phase (Fig. S9b), the CDW moiré superstructure exhibits multiple flat bands and Dirac cones within $\pm 100$ meV of the Fermi energy ($E_{\textrm{F}}$) due to the mixed stacking order (cf. Fig. S6), the structural transformation (cf. Fig. S7), and the electronic reconstruction. Emergent metallicity is evidenced by the density-of-states peak at $E_{\textrm{F}}$ (cf. main text Fig. 4d) as well as the emergent bands. On the $\Gamma$-K-M plane, at 10 meV above $E_{\textrm{F}}$ there is a kagome electronic subsystem that can facilitate in-plane conductivity. On the A-H-L plane, there are several bands crossing $E_{\textrm{F}}$, allowing for non-vertical interlayer hopping. Along the $\Gamma$-A direction, the dispersive bands reflect an out-of-plane metallic character.
Figure S9: Electronic band structures of the CDW moiré superstructure and the 1T-TaS2 C-phase. a, Metallic band structure of the CDW moiré superstructure at different energy windows (blue shading). b, Insulating band structure of the 1T-TaS2 C-phase (i.e., AL-stacking). An energy gap is present from -50 meV to 60 meV around $E_{\textrm{F}}$ (red shading). (Middle) The 1st Brillouin zone of the CDW moiré superstructure. Red dots represent high-symmetry k-points. The aspect ratio of the Brillouin zone ($\approx 0.3$) is adjusted for better visualization of the sampling paths ($\Gamma$-K-M-$\Gamma$-A-H-L-A).
Supplementary References